what does x equal and what does y equal?
x+2y=-5
4x-3y=2

Answer by sdmmadam@yahoo.com(530)

what does x equal and what does y equal?
x+2y=-5
4x-3y=2
x+2y=-5 ----(1) this implies x =( -5-2y) ----(*)
Using (*) in 4x-3y=2 ----(2) that is substituting for x in (2)
4(-5-2y)-3y = 2
-20-8y-3y =2
-20-11y = 2
-20-2 = 11y
-22 = 11y
y = (-22)/11=-2
y=-2 in (*),then x = (-5-2y) = -5-2X(-2) =-5+4 = -1
Answer: x = -1 and y = -2
Verification: Putting x = -1 and y = -2 in x+2y=-5 ----(1)
LHS = x+2y =(-1)+2X(-2) = -1-4 = -5 =RHS
Therefore our values are correct.